# Approximate solutions to the shrinking core model

### v1

@@ 6 Conclusions
In this work, we derived and analyzed the Shrinking Core Model (SCM), which provides a mathematical framework for describing the evolution of a spherical solid particle undergoing reaction with a surrounding fluid. A careful nondimensionalization of the governing equations reveals that the widely used pseudo-steady-state approximation must be applied with caution, since it is generally valid only for solid-gas reactions, where the gas density is significantly lower than that of the solid. For solid-liquid systems, where the fluid and solid densities are comparable, the nondimensional analysis indicates that transient effects may become significant. As a result, applying the pseudo-steady-state approximation in such cases may yield inaccurate predictions and unreliable parameter estimates when fitting experimental data. In addition, the nondimensionalization introduces a key parameter: the ratio of diffusion to reaction time scales. We show that the commonly cited reaction limited and diffusion limited regimes naturally emerge as asymptotic limits of this parameter tending to zero and infinity, respectively, within the pseudo-steady-state framework.
@@
The system’s behavior at early times is also analyzed analytically, and an explicit expression for the small-time solution is derived. This enables the design of a simple fitting procedure that combines the small-time solution with either the pseudo-steady-state or perturbation approximation. Using synthetic data, we demonstrate how the method can be applied to estimate physical parameters from experimental observations with minimal computational cost, offering a practical alternative to full numerical simulations for inferring unknown process parameters.
@@ Acknowledgments
This publication is part of the research project TED2021-131455A-I00 of the Spanish Agencia Estatal de Investigación (AEI) of Ministerio de Ciencia, Innovación y Universidades (MCIU), funding all authors. F. Font, T. Myers and R. Olwande, are also funded by PID2023-146332OB-C21 from AEI of MCIU. F. Font gratefully acknowledges the SGR programme (Ref. 2021-SGR-01045) and the Serra-Hunter Programme of the Generalitat de Catalunya.

### v2

@@ 6 Conclusions
In this work, we derived and analyzed the shrinking core model (SCM), which provides a mathematical framework for describing the evolution of a spherical solid particle undergoing reaction with a surrounding fluid. A careful non-dimensionalisation of the governing equations reveals that the widely used pseudo-steady-state approximation must be applied with caution, since it is generally valid only for solid-gas reactions, where the gas density is significantly lower than that of the solid. For solid-liquid systems, where the fluid and solid densities are comparable, the non-dimensional analysis indicates that transient effects may become significant. As a result, applying the pseudo-steady-state approximation in such cases may yield inaccurate predictions and unreliable parameter estimates when fitting experimental data. In addition, the non-dimensionalisation introduces a key parameter: the ratio of diffusion to reaction time scales. We show that the commonly cited kinetic limited and diffusion limited regimes naturally emerge as asymptotic limits of this parameter tending to zero and infinity, respectively, within the pseudo-steady-state framework.
@@
The system’s behaviour at early times is also analyzed analytically, and an explicit expression for the small-time solution is derived. This enables the design of a simple fitting procedure that combines the small-time solution with either the pseudo-steady-state or perturbation approximation. Using synthetic data, we demonstrate how the method can be applied to estimate physical parameters from experimental observations with minimal computational cost, offering a practical alternative to full numerical simulations for inferring unknown process parameters.
@@ Acknowledgments
This publication is part of the research project TED2021-131455A-I00 of the Spanish Agencia Estatal de Investigación (AEI) of Ministerio de Ciencia, Innovación y Universidades (MCIU), funding all authors. F. Font, T. Myers and R. Olwande, are also funded by PID2023-146332OB-C21 from AEI of MCIU. C. Moreno-Pulido is supported by PID2023-146332OB-C22 from AEI of MCIU. F. Font gratefully acknowledges the SGR programme (Ref. 2021-SGR-01045) and the Serra-Hunter Programme of the Generalitat de Catalunya.